OpenStudy (anonymous):

a salesman earns $200 weekly plus 5% commission on his weekly sales. if he wants to make at least $740 in a week,how much must his sales be?

OpenStudy (anonymous):

740-200=540(he needs to earn this much) so is he sells $515 worth of sales plus 5% commission he will make 740.. 515X5%=25.75(commission) 515+25.75=540.75 540.75+200=740.75

OpenStudy (anonymous):

so his sales must be %515!